# Video-Language Critic: Transferable Reward Functions for Language-Conditioned Robotics

### v1

@@ Abstract
Natural language is often the easiest and most convenient modality for humans to specify tasks for robots. However, learning to ground language to behavior typically requires impractical amounts of diverse, language-annotated demonstrations collected on each target robot. In this work, we aim to separate the problem of what to accomplish from how to accomplish it, as the former can benefit from substantial amounts of external observation-only data, and only the latter depends on a specific robot embodiment. To this end, we propose Video-Language Critic, a reward model that can be trained on readily available cross-embodiment data using contrastive learning and a temporal ranking objective, and use it to score behavior traces from a separate reinforcement learning actor. When trained on Open X-Embodiment data, our reward model enables 2x more sample-efficient policy training on Meta-World tasks than a sparse reward only, despite a significant domain gap. Using in-domain data but in a challenging task generalization setting on Meta-World, we further demonstrate more sample-efficient training than is possible with prior language-conditioned reward models that are either trained with binary classification, use static images, or do not leverage the temporal information present in video data. [^1] Source code, data and trained models are available at https://github.com/minttusofia/video_language_critic.
@@ 1 Introduction
Advances in natural language processing and vision-language representations have enabled a significant increase in the scalability and generalization abilities of learned control policies for robotics. Methods involving large architectures, such as Transformers [42], and internet-scale pretraining have transferred well to both high-level [17, 43] and low-level [5, 21, 34] robotic control. Natural language has many desirable features as a modality for specifying tasks. Unlike structured, hand-designed task sets, natural language is unrestricted and open-domain. Moreover, prompts can be specified as precisely or vaguely as appropriate. While goal images, demonstration videos, or goal states more broadly, have been considered as an alternative open-domain task definition modality [7, 2, 23], they typically have to specify irrelevant environment details, such as the background. Furthermore, language readily supports task definitions with novel combinations of actions, objects and their attributes (e.g. pick up the large green ball) as well as subtask sequencing (e.g. place the toy inside the box, then close the box) in a way that facilitates the policy’s understanding of unseen tasks.
The majority of prior work has proposed to learn language-conditioned policies end-to-end, i.e., directly predicting an action in the robot’s action space given the current state and task description. However, this has several downsides: first, fitting large models on the full problem requires a significant amount of high-quality demonstration data from the target robot domain. Second, the resulting policy depends entirely on the specific robot instance, observation space, and controller type (e.g., joint or task space control) and does not easily transfer to other settings. Moreover, much of the prior work addresses vision-language grounding in robotics purely with imitation learning [5, 21, 34, 27], without attempting to discriminate between low-quality and expert demonstrations. As a result, the resulting policies are inherently limited by the skills of the demonstrators, and no novel solutions can be discovered through planning or interaction. This line of work overlooks performance gains that could be obtained by converting the language prompts to a scalar reward function. Manually defining a well-specified dense reward function to communicate task success is typically laborious and error-prone, and must be repeated for each task. To make progress towards a general-purpose robotic system that can learn human-level skills both in terms of quality (dexterity, robustness) and capability (variety of skills), we argue these systems will need to be able to critique their own behavior, by learning reward functions at scale.
@@
Figure 1: Overview of Vision-Language Critic: Our similarity function, $S_{\theta}$, is trained using video-caption pairs $(v_{1:T}^{i},c^{i})$. The visual encoder (ViT, [9]) is applied separately to each video frame $v_{t}$ to produce a sequence of image features, which are appended to the caption embedding produced by the text encoder. The temporal aggregator then predicts a similarity score for each time step $t$ of the video. We use cross-entropy and sequence ranking objectives to encourage the predicted scores to be high for matching video-caption pairs and low for mismatching pairs, and to monotonously increase over a successful execution.
Recent work in defining language-conditioned rewards for robotic manipulation has used either binary classification [33, 35, 25], contrastive vision-language alignment [38, 22, 26] or reconstruction [15] objectives. However, prior methods have not fully leveraged temporal ordering of frames to encourage increasing scores over a successful episode. We focus on providing a dense reward function that evaluates the actor’s progress throughout an episode by posing learning as a contrastive ranking task.
The advantages and contributions of our approach are as follows:
@@
  • VLC generalizes to unseen tasks through large-scale pretraining and language conditioning and leads to faster policy training than prior reward learning methods.
@@ 2 Related Work
Many prior works have aimed to connect language instructions and vision-based observations in robotics [5, 21, 34, 20, 12] and in video games [10], mostly through large-scale demonstrations [5, 21, 20, 10] or pretraining [34, 12]. However, the majority of approaches have considered imitation-based objectives only, without ranking existing trajectories using offline reinforcement learning (RL) or attempting to outperform prior data using model-based or model-free RL. This has the obvious downside of requiring large amounts of high-quality data from the target domain with low-quality examples already filtered out, and limits the performance of the agent to predicting the mean of the demonstrated behavior instead of improving over it. We instead propose to learn a state-value function from cross-domain offline behavior, which can be optimized using either online, offline or model-based policy training.
@@
Our work is closely related to video retrieval as we seek to move beyond image-language correspondence and match task descriptions with history-aware state representations. As the task of learning representations across time is computationally expensive, many prior works have proposed to start from pretrained, CLIP-like static image features and aggregate them over time, while fine-tuning the aggregation function’s weights on video retrieval [19, 3, 18]. Unlike in video retrieval, we aim to not only assign high alignment scores to full videos, but provide smoothly increasing reward over the whole video to indicate task progress.
@@ 3 Video-Language Critic
3.1 Contrastive video-language training
@@
In order to ensure the reward model learns to discriminate videos based on task completion rather than simply the presence of relevant objects in the scene, it may be beneficial to include failure examples featuring similar environments. Task failures are typically easier to generate than success examples and, as a result, are fairly inexpensive to collect. When available, we leave these videos uncaptioned and treat them only as additional negatives in contrastive learning, and do not include them in the ranking loss.
@@ 4 Experiments
We demonstrate the accuracy and effectiveness of the learned video-language rewards with experimental evaluation on simulated robotic manipulation tasks from the Meta-World benchmark [46]. We evaluate VLC’s ability to inform successful policy training in three settings of increasing difficulty. First, we assess the ability of our model to jointly represent several robotic tasks with a single language-conditioned prediction network in Section 4.1. Second, we test our models’ ability to generalize to unseen Meta-World tasks with the help of vision-language pretraining as well as extrapolation from training tasks in Section 4.2. In Section 4.3, we demonstrate our method’s out-of-domain transfer ability: VLC is used to learn an embodiment-agnostic reward function for any language-conditioned manipulation task by observing a variety of robot actors from Open X-Embodiment [27], a large dataset collected from a variety of real-world robots in different environments. We further report comparisons to prior work, both quantitatively and qualitatively, in Section 4.4. Finally, we demonstrate VLC’s effectiveness in planning with a known dynamical model in Section 4.5.
Reward model training details
We perform hyperparameter selection, ablation studies, and finalize all training details on data from VLMBench [49], a second robotic benchmark. This avoids Meta-World specific tuning and allows for a fairer comparison with prior work. Details on this dataset and our model ablations are included in Appendix A. We observe the same hyperparameters perform well both on VLMBench videos and in interactive RL policy training on Meta-World, which highlights the method’s applicability across domains. Further training details are reported in Appendix B.
Policy training
To train control policies, we use Soft Actor-Critic (SAC) [13], an off-policy RL algorithm. The policy $\pi(a_{t}|s_{t})$ is trained to produce actions $a_{t}$ conditioned on the current environment state $s_{t}$ in order to maximize $\mathbb{E}[R_{t}]$, the expectation of discounted cumulative reward from the current time $t$ onwards: $R_{t}=\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}$, where $r_{t^{\prime}}$ is the reward obtained for taking action $a_{t^{\prime}}$ in state $s_{t^{\prime}}$, $\gamma$ is a discount factor in $[0,1)$ and $T$ is the maximum episode length. The reward function $r$ in our case is either a sparse task completion reward, equal to $1$ if the task has been solved and 00 otherwise, or a learned dense function consisting of a weighted sum of the trained reward function and the sparse reward. As the similarity $S_{\theta}(v_{1:t},c)$ predicted by VLC can take arbitrary values, we apply a few normalization operations to aid in the stability of RL optimization [41]. As predictions for the starting state can vary across initial states, even for the same task, we shift the rewards of the episode such that $r_{1}=0$, as the subsequent behavior should be scored relative to this state. In addition to this, we apply the standard normalization logic from Gymnasium [40], which scales reward values such that their exponential moving average has fixed variance $(1-\gamma)^{2}$.
@@
We evaluate our method in RL policy training on robotic manipulation tasks from the Meta-World benchmark. As the focus of this work is the problem of learning a foundation reward model, we keep to a standard single-task policy training setup and condition on full state information as defined by Meta-World (see Appendix D for more details on the environment definitions). Compared to using image-only observations, this reduces training time and computational cost while allowing us to demonstrate the effectiveness of our reward function. We consider the subset of tasks that can be reliably solved using the dense Meta-World rewards, which are manually specified for each environment based on the full state of the environment. Specifically, we include tasks that can be solved with $\geq$98success within a maximum training length of 800,000 steps. We further split these in half into 12 easy (learned in $<$240,000 steps) and 13 hard tasks (240k – 800k steps).
4.1 Multi-task reward function
The policy training results are shown in Table 4.1, with learning curves for the hard task set in Fig. 2. To summarize learning speed with a single number, we report success rate of the policy evaluated at the training length at which the manually specified Meta-World reward solves the task to $\geq$98success. VLC trained on MW50 enables improved sample efficiency relative to the sparse reward only, which demonstrates that VLC can sufficiently generalize to trajectories not seen in demonstration data, and can effectively represent task progress for multiple tasks at once. However, a few tasks, such as Handle Press, are learned in so few trials even with sparse reward alone that there is little room for improvement in reward design, and learning is instead bottlenecked by the policy training’s sample efficiency. This is why the biggest gains are obtained for the harder tasks.
@@ 4.2 Task generalization to unseen environments
Having verified VLC can effectively represent several tasks using language conditioning, we further evaluate its ability to use language to generalize to entirely unseen tasks. The goal is to see whether knowledge of training tasks, which in part contain similar actions or objects as the held-out tasks (such as Close Faucet in the training set and Open Faucet in the test set), as well as pretrained CLIP language features, are sufficient to allow VLC to generalize to novel tasks. For this purpose, we split the Meta-World tasks into 40 training and 10 test tasks: every 5th task in alphabetical order is reserved for testing. This leaves roughly 1600 successful and 1300 unsuccessful videos as training data – we refer to this subset as MW40. Of the test tasks, 2 are in the easy set, 4 in hard and 4 are unsolved even with the curated single-task Meta-World rewards, and hence not our primary evaluation target.
Success rates of RL policies on these 6 held-out tasks are shown in Table 4.3, with learning curves in Fig. 3. In most tasks, the addition of VLC rewards improves sample efficiency of RL training despite the tasks having never been seen in reward function training, with average success rates 34 percentage points over the sparse baseline and an average sample efficiency improvement of over 5x.
4.3 Embodiment generalisation to unseen domains
We successfully train policies (see Table 4.1 and Fig. 2) using Open X trained models despite a significant domain gap, highlighting the generalizability of large-scale vision-language training. We obtain an average 2.1x sample efficiency gain relative to the sparse reward in tasks that are solved by both rewards (sample efficiency is ill-defined if either does not solve the task), with particularly large improvements in Handle Pull Side (7x), Reach Wall (7x) and Slide Plate (6x) and an average 5 percentage point success rate increase across all 25 tasks despite misrepresenting a few tasks. Note that unlike RT-X [27], our method does not use action labels, and remains equally applicable on observation-only data.
@@ 4.3 Embodiment generalisation to unseen domains
easy (2) 45 $\pm$ 16 47 $\pm$ 26 30 $\pm$ 10 18 $\pm$ 10 22 $\pm$ 11 51 $\pm$ 15 39 $\pm$ 12
hard (4) 72 $\pm$ 19 57 $\pm$ 12 63 $\pm$ 15 35 $\pm$ 11 24 $\pm$ 14 87 $\pm$ 17 42 $\pm$ 10
mean 63 $\pm$ 18 53 $\pm$ 13 39 $\pm$ 11 29 $\pm$ 19 16 $\pm$ 10 75 $\pm$ 17 41 $\pm$ 18
@@ 4.4 Comparison to prior work
To validate the benefits of VLC, we compare its performance to prior language-conditioned reward models LOReL [25], RoboCLIP [38], LIV [22], Voltron [15] and R3M [26], each fine-tuned on MW40. Training and implementation details are included in Appendix C.
We find VLC’s combination of cross-entropy and the sequential ranking objective as well as full video conditioning to produce more informative reward predictions than existing methods, as shown by faster policy training on average in Table 4.3 and Fig. 3. Moreover, on qualitative inspection of the shape of the predicted rewards (Fig. 4), we find VLC’s outputs to better distinguish successes from failures compared to either LOReL or RoboCLIP, the two strongest baselines. Thanks to its broad coverage of execution history and the sequential loss term, VLC produces rewards that more smoothly increase over time for successful executions than either prior method.
@@ 4.5 Model-based evaluation
As a pretrained reward function, VLC can also be very useful in the context of model-based planning: any imagined future state can be scored, allowing the planner to identify the best action plan. We demonstrate this in a proof-of-concept experiment, where we do not learn the model but instead assume access to a known transition model as well as action primitives. The action primitives include grasping and reaching, parameterized by target positions (such as the locations of objects detected in the scene), and are defined using segments from the expert policies available in Meta-World. We evaluate VLC’s ability to identify the action primitive with the correct execution for a held-out task.
In each task, we compare one successful trajectory with 5 unsuccessful ones with randomly sampled target positions. Two action primitives are required to solve Assembly and Pick Out of Hole: first grasping the required object, then moving it as specified, such as lifting it out of the hole. The action primitive is selected with $\arg\max_{a}S_{\theta}(o(s_{t+1}),c)$ where $a$ is the action primitive that transitions state $s_{t}$ to $s_{t+1}$, and $o$ is a function describing the observation visible to $S_{\theta}$, in this case a video.
In Table 3, we report the accuracy of our model in identifying executions solving each held-out task, averaged over 50 episodes. VLC generalizes well to these tasks zero-shot, with 80% mean accuracy.
Table 3: Accuracy (%) of VLC-MW40 in identifying a successful execution out of 6 samples.
100 20 96 100 40 94 54 82 100 100 97 100 80
5 Limitations
To learn reward functions from cross-embodiment data, a shared input representation, typically images or video, is required. This limits the application to tasks whose progress can be evaluated from vision; however, additional robot-specific variables could be added to the policy’s input representation. Furthermore, the size of the dataset we use in cross-embodiment experiments is significant and imposes considerable computational cost, which is why we chose to focus on Meta-World task generalization when comparing to prior work, in order to be able to adequately train several methods. Further work is needed to investigate the performance with smaller datasets. For tasks that clearly deviate from the Open X training tasks, such as in their length or specificity, additional data collection might be beneficial. In our formulation, we focused on tasks whose rewards can be simplified to goal reaching, but other rewards may be relevant (such as penalizing large actions). We further assumed that a sparse task completion reward can be provided, which may not always be the case.
In this work, we focused on simulated robotic manipulation tasks, and evaluation on a real robot is left for further work. Moreover, extending VLC to learn from videos of humans performing manipulation tasks, such as from the Something-Something dataset [11], is a promising avenue for pushing the generalization capabilities of pretrained reward functions in the future.

### v2

@@ Abstract
Natural language is often the easiest and most convenient modality for humans to specify tasks for robots. However, learning to ground language to behavior typically requires impractical amounts of diverse, language-annotated demonstrations collected on each target robot. In this work, we aim to separate the problem of what to accomplish from how to accomplish it, as the former can benefit from substantial amounts of external observation-only data, and only the latter depends on a specific robot embodiment. To this end, we propose Video-Language Critic, a reward model that can be trained on readily available cross-embodiment data using contrastive learning and a temporal ranking objective, and use it to score behavior traces from a separate actor. When trained on Open X-Embodiment data, our reward model enables 2x more sample-efficient policy training on Meta-World tasks than a sparse reward only, despite a significant domain gap. Using in-domain data but in a challenging task generalization setting on Meta-World, we further demonstrate more sample-efficient training than is possible with prior language-conditioned reward models that are either trained with binary classification, use static images, or do not leverage the temporal information present in video data.
Keywords: Vision-Language Models, Reinforcement Learning, Reward Learning
@@ 1 Introduction
Advances in natural language processing and vision-language representations have enabled a significant increase in the scalability and generalization abilities of learned control policies for robotics. Methods involving large architectures, such as Transformers [1], and internet-scale pretraining have transferred well to both high-level [2, 3] and low-level [4, 5, 6] robotic control. Natural language has many desirable features as a modality for specifying tasks. Unlike structured, hand-designed task sets, natural language is unrestricted and open-domain. Moreover, prompts can be specified as precisely or vaguely as appropriate. While goal images, demonstration videos, or goal states more broadly, have been considered as an alternative open-domain task definition modality [7, 8, 9], they typically have to specify irrelevant environment details, such as the background. Furthermore, language readily supports task definitions with novel combinations of actions, objects and their attributes, as well as subtask sequencing, in a way that facilitates the policy’s understanding of unseen tasks.
Most prior work has proposed to learn language-conditioned policies end-to-end, i.e., directly predicting an action in the robot’s action space given the current state and task description. However, this has several downsides: first, fitting large models on the full problem requires a significant amount of high-quality demonstration data from the target robot domain. Second, the resulting policy depends entirely on the specific robot instance, observation space, and controller type (e.g., joint or task space control) and does not easily transfer to other settings. Moreover, much of the prior work addresses vision-language grounding in robotics purely with imitation learning [4, 5, 6, 10], without attempting to discriminate between low-quality and expert demonstrations. As a result, the resulting policies are inherently limited by the skills of the demonstrators, and no novel solutions can be discovered through planning or interaction. This line of work overlooks performance gains that could be obtained by converting the language prompts to a scalar reward function. Manually defining a well-specified dense reward function to communicate task success is typically laborious and error-prone, and must be repeated for each task. To make progress towards a general-purpose robotic system that can learn human-level skills both in terms of quality (dexterity, robustness) and variety of skills, we argue these systems will need to be able to critique their own behavior, by learning reward functions at scale.
@@
Figure 1: Overview: Our similarity function, $S_{\theta}$, is trained using video-caption pairs $(v_{1:T}^{i},c^{i})$. The visual encoder (ViT, [12]) is applied separately to each video frame $v_{t}$ to produce a sequence of image features, which are appended to the caption embedding produced by the text encoder. The temporal aggregator then predicts a similarity score for each time step $t$ of the video. We use cross-entropy and sequence ranking objectives to encourage the predicted scores to be high for matching video-caption pairs and low for mismatching pairs, and to monotonously increase over a successful execution.
Recent work in language-conditioned rewards for robotics has used either binary classification [13, 14, 15], contrastive vision-language alignment [16, 17, 18] or reconstruction [19] objectives. However, they have not fully leveraged temporal ordering of frames to encourage increasing scores over successful episodes. In contrast, VLC provides a dense reward evaluating in-episode progress, learned through contrastive ranking. The advantages and contributions of our approach are as follows:
@@
  • VLC generalizes to unseen tasks through large-scale pretraining and language conditioning and leads to faster policy training than 5 prior reward learning methods.
@@ 2 Related Work
Many prior works have aimed to connect language instructions and vision-based observations in robotics [4, 5, 6, 20, 21] and in video games [22], mostly through large-scale demonstrations [4, 5, 20, 22] or pretraining [6, 21]. However, the majority of approaches have considered imitation-based objectives only, without ranking existing trajectories or attempting to outperform prior data. We instead propose to learn a state-value function from cross-domain offline behavior, which can be optimized using either online, offline or model-based policy training.
@@
Our work is related to video retrieval as we seek to move beyond image-language correspondence and match task descriptions with history-aware state representations. As learning representations across time is computationally expensive, many prior works have proposed to start from pretrained image-language representations and aggregate them over time, while fine-tuning the aggregation function’s weights on video retrieval [26, 27, 28]. Unlike in video retrieval, we aim to not only assign high alignment scores to full videos, but provide smoothly increasing reward over the whole video to indicate task progress.
@@ 3 Video-Language Critic
Contrastive video-language training
@@
In order to ensure the reward model learns to discriminate videos based on task completion rather than simply the presence of relevant objects in the scene, it may be beneficial to include failure examples featuring similar environments. Task failures are typically easier to generate than success examples and thus fairly inexpensive to collect. When available, we leave these videos uncaptioned and treat them only as additional negatives in contrastive learning, and do not include them in the ranking loss.
@@ 4 Experiments
We evaluate the accuracy and effectiveness of the learned video-language rewards on simulated robotic manipulation tasks from the Meta-World benchmark [11]. We evaluate VLC’s ability to inform successful policy training in three settings of increasing difficulty. First, we assess the ability of our model to jointly represent several robotic tasks with a single language-conditioned prediction network in Section 4.1. Second, we test our models’ ability to generalize to unseen Meta-World tasks with the help of vision-language pretraining as well as extrapolation from training tasks in Section 4.2. In Section 4.3, we demonstrate our method’s out-of-domain transfer ability: VLC is used to learn an embodiment-agnostic reward function for any language-conditioned manipulation task by observing a variety of robot actors from Open X-Embodiment [10], a large dataset collected from a variety of real-world robots in different environments. We further report comparisons to prior work, both quantitatively and qualitatively, in Section 4.4, and demonstrate VLC’s effectiveness in planning with a known dynamical model in Section 4.5. Finally, in Appendix E, we investigate the model’s robustness to visual variations (brightness shift and image noise).
Training details
We perform hyperparameter selection, ablation studies, and finalize all training details on data from VLMBench [40], a second robotic benchmark. This avoids Meta-World specific tuning and allows for a fairer comparison with prior work. Details on this dataset and our model ablations are included in Appendix A. We observe the same hyperparameters perform well both on VLMBench videos and in interactive RL policy training on Meta-World, which highlights the method’s applicability across domains. Further VLC training details are reported in Appendix B. To train control policies, we use Soft Actor-Critic (SAC) [41], an off-policy RL algorithm. As the reward function, we use either a sparse task completion reward, equal to $1$ if the task has been solved and 00 otherwise, or a weighted sum of the learned VLC reward and the sparse reward. As the similarity $S_{\theta}(v_{1:t},c)$ predicted by VLC can take arbitrary values, we apply a few normalization operations to aid in the stability of RL optimization [42] (details in Appendix B.2).
@@
We evaluate VLC in RL policy training on robotic manipulation tasks from the Meta-World benchmark. As the focus of this work is the problem of learning a foundation reward model, we keep to a standard single-task policy training setup and condition on full state information as defined by Meta-World (see Appendix C for more details on the environment definitions). Compared to using image-only observations, this reduces training time and computational cost while allowing us to demonstrate the effectiveness of our reward function. We consider the subset of tasks that can be reliably solved using the dense Meta-World rewards, which are manually specified for each environment based on the full state of the environment. Specifically, we include tasks that can be solved with $\geq$98success within a maximum training length of 800,000 steps. We further split these in half into 12 easy (learned in $<$240,000 steps) and 13 hard tasks (240k – 800k steps).
@@
4.1 Multi-task reward function
The policy training results are shown in Table 4, with learning curves for the hard task set in Fig. 2. To summarize learning speed with a single number, we report success rate of the policy evaluated at the training length at which the manually specified Meta-World reward solves the task to $\geq$98success. VLC trained on MW50 enables improved sample efficiency relative to the sparse reward only, which demonstrates that VLC can sufficiently generalize to trajectories not seen in demonstration data, and can effectively represent task progress for multiple tasks at once. However, a few tasks, such as Handle Press, are learned in so few trials even with sparse reward alone that there is little room for improvement in reward design, and learning is instead bottlenecked by the policy training’s sample efficiency. This is why the biggest gains are obtained for the harder tasks.
@@ 4.2 Task generalization to unseen environments
Next, we evaluate VLC’s ability to generalize to entirely unseen tasks using language conditioning. For this purpose, we split Meta-World into 40 training and 10 test tasks (every 5th task alphabetically). This leaves roughly 1600 successful and 1300 unsuccessful videos as training data – we refer to this subset as MW40. Of the test tasks, 2 are in the easy set, 4 in hard and 4 are unsolved even with the curated single-task Meta-World rewards, and hence not our primary evaluation target.
Success rates of RL policies on these 6 held-out tasks are shown in Table 4.2 (calculated similarly to Table 4), with learning curves in Fig. 3. In most tasks, the addition of VLC rewards improves sample efficiency of RL training despite the tasks having never been seen in reward function training, with average success rates 34 percentage points over the sparse baseline and an average sample efficiency improvement of over 5x.
In addition to task generalization, we further evaluate VLC in out-of-distribution visual conditions in Appendix E.
@@
easy (2) 57 $\pm$ 21 46 $\pm$ 26 30 $\pm$ 10 18 $\pm$ 10 22 $\pm$ 11 58 $\pm$ 22 39 $\pm$ 12
hard (4) 60 $\pm$ 16 51 $\pm$ 11 63 $\pm$ 15 35 $\pm$ 11 24 $\pm$ 14 79 $\pm$ 18 42 $\pm$ 10
mean 59 $\pm$ 19 50 $\pm$ 12 39 $\pm$ 11 29 $\pm$ 19 16 $\pm$ 10 72 $\pm$ 11 41 $\pm$ 18
4.3 Embodiment generalization to unseen domains
We successfully train policies (see Table 4 and Fig. 2) using Open X trained models despite a significant domain gap, highlighting the generalizability of large-scale vision-language training. We obtain an average 2.1x sample efficiency gain over the sparse reward in tasks solved by both rewards (sample efficiency is ill-defined if either does not solve the task), with particularly large improvements in Handle Pull Side (7x), Reach Wall (7x) and Slide Plate (6x) and an average 5 percentage point success rate increase across all 25 tasks despite misrepresenting a few tasks. Note that unlike RT-X [10], our method does not use action labels, and remains equally applicable on observation-only data.
@@ 4.4 Comparison to prior work
To validate the benefits of VLC, we compare its performance to prior language-conditioned reward models LOReL [15], RoboCLIP [16], LIV [17], Voltron [19] and R3M [18], each fine-tuned on MW40. Training and implementation details are included in Appendix D.
We find VLC’s combination of cross-entropy and the sequential ranking objective as well as full video conditioning to produce more informative reward predictions than existing methods, as shown by faster policy training on average in Table 4.2 and Fig. 3. For increased statistical significance, we use 5 additional random seeds (10 total) for VLC and the two strongest baselines: RoboCLIP and LOReL. Moreover, on qualitative inspection of the shape of the predicted rewards (Appendix F), we find VLC’s outputs to better distinguish successes from failures compared to either RoboCLIP or LOReL. Thanks to its broad coverage of execution history and the sequential loss term, VLC also produces rewards that more smoothly increase over successful episodes than either prior method.
@@ 4.5 Model-based evaluation
As a pretrained reward function, VLC can also inform model-based planning. We demonstrate this in a proof-of-concept experiment, where we do not learn the model but instead assume access to a known transition model as well as action primitives. The action primitives include grasping and reaching, parameterized by target positions (such as the locations of objects detected in the scene), and are defined using segments from the expert policies available in Meta-World. We evaluate VLC’s ability to identify the action primitive with the correct execution for a held-out task by assigning it a higher score than for incorrect executions. In each task, we compare one successful trajectory with 5 unsuccessful ones with randomly sampled target positions.
We find VLC to generalize well to these tasks zero-shot, with 80% mean accuracy on held-out tasks, outperforming RoboCLIP (46%) and LOReL (36%). A breakdown per task is shown in Appendix G.